OpenStudy (anonymous):

Match the sequence with the correct indicated term. a43 for 5,9,13,...

OpenStudy (anonymous):

here are the answer choices 340 -47 25/2 -175 173 -25/2

OpenStudy (anonymous):

Here \(a_1=5\), \(a_2=9\), \(a_3=13\) Use \(T_n=a_1+(n-1)d\) to find d. \(T_2=5+(2-1)d\) \(9=5+d\) \(d=9-5\) \(d=4\) Again use \(T_n=a_1+(n-1)d\) to find \(a_{43}\) \(T_{43}=5+(43-1)4\) \(=5+42*4\) \(=5+168\) \(=173\) Is that clear?

OpenStudy (anonymous):

im sorry i was doing something give me time to look over it please

OpenStudy (anonymous):

Ha k.

OpenStudy (anonymous):

where did the d come from and the Tn?

OpenStudy (anonymous):

It is a general formula. \(T_n=a_n\) d=common difference. Nw is it clear?

OpenStudy (anonymous):

yes i guess

OpenStudy (anonymous):

idk im kinda lost is there another way to explain it?

OpenStudy (anonymous):

can we just move to another one

OpenStudy (anonymous):

i have other examples

OpenStudy (anonymous):

\(Tn=a_1+(n-1)d\).is a general formula Here \(T_n\)=\(n^{th}\) term \(a_1\)=First term d=common difference. Ya sure post them.

OpenStudy (anonymous):

ohk a1 =20, d=4, n=81

OpenStudy (anonymous):

Vat do u need to find here?

OpenStudy (anonymous):

the same question

OpenStudy (anonymous):

same answer choices

OpenStudy (anonymous):

I mean whether u need to find the sum or \(a_8\)?

OpenStudy (anonymous):

its ohk lol i kinda understand now tht i keep lookin at it im going to just follow your steps